# Classification of thin Jordan schemes

### v1

@@ 2.5. Algebraic fusions of J-configurations
Given a J-configuration ${\mathfrak{S}}=(\Omega,S)$, a permutation $\alpha\in{\sf Sym}(S)$ is called an algebraic J-automorphism of $\mathfrak{S}$} if its linear extension to $\mathbb{C}[S]$ is an automorphism of the Jordan algebra $(\mathbb{C}[S],\star)$. All algebraic J-automorphisms of $\mathfrak{S}$} form a subgroup of ${\sf Sym}(S)$ that will be denoted by $\mathsf{JAut}({\mathfrak{S}})$. Notice that the transpose permutation $\tau:a\mapsto a^{t}$ is always an algebraic J-automorphism of $(\mathbb{C}[S],\star)$. The group $\mathsf{JAut}({\mathfrak{S}})$ consists of all permutations of ${\sf Sym}(S)$ which preserve the tensor of structure constants of Jordan multiplication with respect to the standard basis. In other words, a permutation $u\mapsto u^{\prime}$ of $S$ belongs to $\mathsf{JAut}({\mathfrak{S}})$ if and only if
$\forall_{u,v,w\in S}:\ p_{\{u^{\prime},v^{\prime}\}}^{w^{\prime}}=p_{\{u,v\}}^{w}$
The following statement collects some properties of subgroups of $\mathsf{JAut}({\mathfrak{S}})$.
@@
The following properties hold.
@@
    The relations $\Phi(s),s\in S$ form a JC on $\Omega$.
@@
Proof of part (a). One can easily check that for any pair $r,s\in S$, it holds that
@@
hereby, implying $s=r^{t}\iff\phi(s)=\phi(r)^{t}$.
Proof of part (b). It follows from part (a) that
@@
If $s^{t}=\phi(s)$ for some $\phi\in\Phi$, then $\Phi(s^{t})=\Phi(\phi(s))=\Phi(s)$ implying that $\Phi(s)$ is symmetric.
If $s^{t}\neq\phi(s)$ for every $\phi\in\Phi$, then $\psi(s^{t})\neq\phi(s)$, and, consequently, $\psi(s^{t})\cap\phi(s)=\emptyset$ for any pair of $\psi,\phi\in\Phi$. Therefore $\Phi(s^{t})\cap\Phi(s)=\bigcup_{\psi,\phi\in\Phi}(\psi(s^{t})\cap\phi(s))=\emptyset$. Hence $\Phi(s)$ is anti-symmetric in this case.
[…]
@@
Since $\phi(I_{\Omega})=I_{\Omega},\phi(J_{\Omega})=J_{\Omega}$, the subspace $^{\Phi}\!\!{\mathcal{A}}=\{X\in{\mathcal{A}}\mid\forall_{\phi\in\Phi}\ \phi(X)=X\}$ is a coherent Jordan algebra. It is easy to check that the matrices $\{\underline{\Phi(s)}\}_{s\in S}$ form the standard basis of $^{\Phi}\!\!{\mathcal{A}}$. ∎
It follows from the above statement that for any $\Phi\leq\mathsf{JAut}({\mathfrak{S}})$ the pair $^{\Phi}{{\mathfrak{S}}}:=(\Omega,\{\Phi(s)\}_{s\in S})$ is also a JC which will be called an algebraic fusion of $\mathfrak{S}$}.
@@ 2.6. Constructing Jordan configurations from coherent ones
If ${\mathfrak{S}}=(\Omega,S)$ is a CC, then it is also a JC. Recall that a permutation $\phi$ of $S$ is called an algebraic automorphism (anti-automorphism) of $\mathfrak{S}$} if its linear extension to $\mathcal{A}:=\langle\{\underline{s}\}_{s\in S}\rangle$ is an automorphism (anti-automorphism) of $(\mathcal{A},\cdot)$. The group $\mathsf{AAut}({\mathfrak{S}})$ of all algebraic automorphisms of $\mathfrak{S}$} is contained in $\mathsf{JAut}({\mathfrak{S}})$. Also, $\tau\in\mathsf{JAut}({\mathfrak{S}})$. Every anti-automorphism of $\mathfrak{S}$} is a composition of $\tau$ and some $\phi\in\mathsf{AAut}({\mathfrak{S}})$. All automorphisms and anti-automorphisms of $\mathfrak{S}$} form a subgroup of $\mathsf{JAut}({\mathfrak{S}})$ which will be denoted by $\mathsf{TAut}({\mathfrak{S}})$. If $\mathfrak{S}$} is commutative, then $\mathsf{AAut}({\mathfrak{S}})=\mathsf{TAut}({\mathfrak{S}})=\mathsf{JAut}({\mathfrak{S}})$. If $\mathfrak{S}$} is not commutative, then $\tau\not\in\mathsf{AAut}({\mathfrak{S}})$ and $\mathsf{TAut}({\mathfrak{S}}):\mathsf{AAut}({\mathfrak{S}})]=2$. It is easy to check that $\mathsf{TAut}({\mathfrak{S}})=\mathsf{AAut}({\mathfrak{S}})\times\langle\tau\rangle$. It is obvious that $\mathsf{TAut}({\mathfrak{S}})\leq\mathsf{JAut}({\mathfrak{S}})$. In general, the inclusion could be proper. The smallest example known to us has order $16$ and was found by computer search.
If ${\mathfrak{S}}=(\Omega,S)$ is a thin association scheme, then $S$ is a regular subgroup of ${\sf Sym}(\Omega)$. In this case a permutation $\phi\in{\sf Sym}(S)$ belongs to $\mathsf{JAut}({\mathfrak{S}})$ if and only if it satisfies $\phi(st)\in\{\phi(s)\phi(t),\phi(t)\phi(s)\}$ for all $s,t\in S$. It was shown in [13] that any such map is either automorphism or anti-automorphism [^1] The authors are grateful to R. Lipyanski who informed them about the paper [13]. Thus for thin AS we always have $\mathsf{TAut}({\mathfrak{S}})=\mathsf{JAut}({\mathfrak{S}})$.
We call a JC $\mathfrak{S}=(\Omega,S)$ non-autonomous if it is an algebraic fusion of a coherent configuration $\mathfrak{C}=(\Omega,C)$ via a subgroup $\Phi\leq\mathsf{JAut}(\mathfrak{C})$. A JC will be called autonomous if it cannot be obtained as an algebraic fusion of some CC sharing the same point set. Notice that in [12] Jordan schemes were divided into two types: proper and improper. Improper are those which are symmetrizations of ASs. The others are proper. Under this definition, any non-symmetric JS is automatically proper. Regarding symmetric Jordan schemes, there are proper ones which are non-autonomous. The smallest example known to us is of order $24$ (we know that there are no examples of order less than $16$).
@@
Computer calculations show that $\text{AS}(24,526)$, $\text{AS}(24,597)$, $\text{AS}(24,669)$ from the catalog of small association schemes by Hanaki and Miyamoto (see [7], available in GAP [6] via the COCO2P package [11]), each has two algebraic fusions that are proper symmetric Jordan schemes.
It looks like the proper JSs constructed in [12] are also autonomous, but we did not prove this. We can claim this only for the smallest example of order $15$ constructed there.
@@ 3. Thin Jordan schemes
It is well-known that each basic graph of an association scheme $(\Omega,S)$ is regular. Therefore we always have that $|S|\leq|\Omega|$ where the equality holds if and only if each basic relation $s\in S$ is a permutation on $\Omega$. Such a scheme is called thin and it is not difficult to check that an association scheme $(\Omega,S)$ is thin if and only if $S$ is a regular subgroup of ${\sf Sym}(\Omega)$. Since every group has a regular representation, one can consider groups as a particular case of association schemes.
Although symmetric Jordan schemes are always regular, it is no longer true for nonsymmetric ones, as the following simple example shows.
@@
The full matrix algebra $M_{\Omega}({\mathbb{F}})$ is a discrete coherent algebra of rank $|\Omega|^{2}$. In the case of $\Omega=\{1,2\}$ the mapping $\phi:\begin{bmatrix}a&b\\ c&d\\ \end{bmatrix}\rightarrow\begin{bmatrix}d&b\\ c&a\\ \end{bmatrix}$ is an anti-automorphism. Therefore, the subspace ${\mathcal{A}}:={}^{\langle\phi\rangle}M_{\{1,2\}}({\mathbb{F}})=\left\{\begin{bmatrix}a&b\\ c&a\\ \end{bmatrix}\,\vline\,a,b,c\in{\mathbb{F}}\right\}$ is a Jordan coherent algebra. It has order two and rank three. The corresponding coherent J-configuration $(\Omega=\{1,2\},S=\{s_{0}=I_{\Omega},s_{1}=\{(1,2)\},s_{2}=s_{1}^{t}\})$ is homogeneous but not regular. It’s rank-to-order ratio $|S|/|\Omega|$ is equal to $3/2$.
The following statement shows that this is the maximal value of a rank-to-order ratio for a non-regular JS.
@@
Let $(\Omega,S)$ be a non-regular JS. Then $|S|/|\Omega|\leq 3/2$ where the equality holds if and only if all basic relations are thin.
@@
By [12, Proposition 2.2] and its proof there exists a partition $\Omega=\Omega_{1}\cup\Omega_{0}$ with $|\Omega_{1}|=|\Omega_{0}|$ such that every regular relation is contained in $\Omega_{0}^{2}\cup\Omega_{1}^{2}$ and each non-regular one is contained either in $\Omega_{0}\times\Omega_{1}$ or in $\Omega_{1}\times\Omega_{0}$. Thus $S$ splits into a disjoint union $S=S_{r}\cup S_{0,1}\cup S_{1,0}$ where $S_{r}$ stands for all regular relations while $S_{i,1-i}:=\{s\in S\mid\subseteq\Omega_{i}\times\Omega_{1-i}\}$.
It follows from $\sum_{s\in S_{i,1-i}}|s|=|\Omega_{i}|\cdot|\Omega_{1-i}|$ and $|s|=n_{s}|\Omega_{i}|$ that $\sum_{s\in S_{i,1-i}}n_{s}=|\Omega|/2$. Since all summands in the latter sum are non-negative integers, we conclude $|S_{i,1-i}|\leq|\Omega|/2$ for $i=0,1$.
@@
Arguing as before we conclude that $|S_{r}|\leq|\Omega|/2$. Combining altogether we obtain
@@
It follows from the above arguments that an equality $|S|=\frac{3}{2}|\Omega|$ happens if and only if $n_{s}=1$ for all $s\in S$. ∎
Since the maximal value of rank-order ratio is attained by non-regular thin Jordan schemes, our first goal is to classify all of them. This is done in the next subsection.
3.1. Non-regular thin Jordan schemes
Let us start this section by describing how every commutative homogeneous coherent algebra gives rise to a non-regular homogeneous coherent Jordan algebra.
@@ 3.1. Non-regular thin Jordan schemes
Let $\mathcal{C}\leq M_{n}(\mathbb{F})$ be a commutative homogeneous coherent algebra. Then we define $\mathcal{J}(\mathcal{C})\subseteq M_{2n}(\mathbb{F})$ according to:
@@
Clearly, $\mathcal{J}(\mathcal{C})$ is a linear subspace of $M_{2n}(\mathbb{F})$. Since $\mathcal{C}$ itself is closed under Schur-Hadamard multiplication, so is $\mathcal{J}(\mathcal{C})$. It remains to show the closure under Jordan product: to this end, let $A,B,C,D,E,F\in\mathcal{C}$. We compute
@@
An important special case is when $\mathcal{C}$ is the adjacency algebra of a thin association scheme coming from an abelian group $G$. In this case, to simplify notation, instead $\mathcal{J}(\mathcal{A}(G,G))$ we write $\mathcal{J}(G)$. Clearly, for every abelian group $G$ we have that $\mathcal{J}(G)$ is a non-regular thin homogeneous Jordan algebra.
@@
Let $\mathfrak{S}=(\Omega,\mathcal{C})$ be a non-regular thin Jordan scheme. and let $\mathcal{A}$ be its adjacency algebra. Then there exists an abelian group $G$ and a permutation matrix $P$, such that $\mathcal{A}=P^{-1}\mathcal{J}(G)P$.
@@
Let $\mathfrak{J}=(\Omega,\mathcal{C})$ be a non-regular thin Jordan scheme and let $\mathcal{A}$ be its adjacency algebra. Let ${\mathcal{S}}=\{A_{i}\}_{i=1}^{r}$ be the standard basis of $\mathcal{A}$. By [12, Proposition 2.2], each element of $\mathcal{S}$} is of one of the following shapes:
@@
where $A,B,C,D$ are certain matrices of order $n$, and where $O$ denotes the zero matrix. Since $\mathfrak{S}$ is thin, the matrices $A,B,C,D$ are permutation matrices.
Let
@@
Note that $\mathcal{K}=\mathcal{H}^{t}=\mathcal{H}^{-1}$ and that $\sum\mathcal{G}_{1}=\sum\mathcal{G}_{2}=\sum\mathcal{H}=\sum\mathcal{K}=J$. Moreover, $|\mathcal{G}_{1}|=|\mathcal{G}_{2}|=|\mathcal{H}|=|\mathcal{K}|=n$.
Next we show that:
@@
Indeed, suppose that $\left[\begin{smallmatrix}A&O\\[2.84526pt] O&B_{1}\end{smallmatrix}\right],\left[\begin{smallmatrix}A&O\\[2.84526pt] O&B_{2}\end{smallmatrix}\right]\in{\mathcal{S}}$, but $B_{1}\neq B_{2}$. Then $\left[\begin{smallmatrix}O&O\\[2.84526pt] O&B_{1}-B_{2}\end{smallmatrix}\right]\in\mathcal{A}$. But then also
$\begin{bmatrix}O&O\\[2.84526pt] O&B_{1}-B_{2}\end{bmatrix}\circ\begin{bmatrix}A&O\\[2.84526pt] O&B_{1}\end{bmatrix}=\begin{bmatrix}O&O\\[2.84526pt] O&B_{1}\end{bmatrix}$
As $\mathcal{A}$ is closed under taking powers of matrices, we obtain that $\left[\begin{smallmatrix}O&O\\[2.84526pt] O&I\end{smallmatrix}\right]\in\mathcal{A}$. However, this is a contradiction with the assumption that $\mathfrak{J}$ is a Jordan-scheme. It follows that $B_{1}=B_{2}$.
Thus we created a bijection between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. Let us denote it by $A\mapsto\varphi(A)$. Clearly, $\mathcal{G}_{1}$ is the standard basis of a thin, homogeneous coherent $J$-algebra. The same is true for $\mathcal{G}_{2}$.
Next we show that
@@
To see this, observe that for all $C\in\mathcal{H}$, $D\in\mathcal{K}$ we have
@@
Since $|C\mathcal{K}|=|\mathcal{K}|=|\mathcal{G}_{1}|$, there exists a $D\in\mathcal{K}$ such that $A=CD$. By (1), for this $D$ holds $\varphi(A)=DC$.
Next we show that
@@
To this end let $A\in\mathcal{G}_{1}$. By (2), there exists $D\in\mathcal{K}$, such that $A=CD$ and $\varphi(A)=DC$. From this follows that $D=C^{-1}A$. hence $\varphi(A)=C^{-1}AC$.
Note that conjugation of $\mathcal{A}$ by a permutation matrix $P$ ($X\mapsto PXP^{-1}$) preserves $\cdot$, $\star$, $\circ$, $(-)^{t}$. In other words, the resulting algebra is a homogeneous coherent $J$-algebra $\mathcal{A}^{\prime}=P\mathcal{A}P^{-1}$, and the corresponding Jordan schemes are combinatorially isomorphic.

### v2

@@ 2.5. Algebraic fusions of J-configurations
Given a J-configuration ${\mathfrak{S}}=(\Omega,S)$, a permutation $\alpha\in{\sf Sym}(S)$ is called an algebraic J-automorphism of $\mathfrak{S}$} if its linear extension to $\mathbb{C}[S]$ is an automorphism of the Jordan algebra $(\mathbb{C}[S],\star)$. All algebraic J-automorphisms of $\mathfrak{S}$} form a subgroup of ${\sf Sym}(S)$ that will be denoted by $\mathsf{JAut}({\mathfrak{S}})$. Note that the transpose permutation $\tau:a\mapsto a^{t}$ is always an algebraic J-automorphism of $\mathfrak{S}$}. The group $\mathsf{JAut}({\mathfrak{S}})$ consists of all permutations of ${\sf Sym}(S)$ that preserve the tensor of structure constants of Jordan multiplication with respect to the standard basis. In other words, a permutation $u\mapsto u^{\prime}$ of $S$ belongs to $\mathsf{JAut}({\mathfrak{S}})$ if and only if
$\forall_{u,v,w\in S}:\ p_{\{u^{\prime},v^{\prime}\}}^{w^{\prime}}=p_{\{u,v\}}^{w}.$
The following statement collects certain properties of subgroups of $\mathsf{JAut}({\mathfrak{S}})$.
@@
Each of the following holds:
@@
    The relations $\Phi(s),s\in S$ form a Jordan configuration on $\Omega$.
@@
Proof of (a). One readily checks that for any pair $r,s\in S$,
@@
thereby implying that $s=r^{t}\iff\phi(s)=\phi(r)^{t}$.
Proof of (b). It follows from part (a) that
@@
If $s^{t}=\phi(s)$ for some $\phi\in\Phi$, then $\Phi(s^{t})=\Phi(\phi(s))=\Phi(s)$, which implies $\Phi(s)$ is symmetric. If, on the other hand, $s^{t}\neq\phi(s)$ for every $\phi\in\Phi$, then $\psi(s^{t})\neq\phi(s)$ and consequently $\psi(s^{t})\cap\phi(s)=\emptyset$ for every pair $\psi,\phi\in\Phi$. Therefore $\Phi(s^{t})\cap\Phi(s)=\bigcup_{\psi,\phi\in\Phi}(\psi(s^{t})\cap\phi(s))=\emptyset$, whereby $\Phi(s)$ is anti-symmetric in this case.
Proof of (c). Denote ${\mathcal{A}}:=\mathbb{C}[S]$ and extend each $\phi\in\Phi$ on $\mathcal{A}$} linearly via $\phi(\underline{s}):=\underline{\phi(s)}$. We will use the same notation for the induced linear operator on $\mathcal{A}$}. Then
@@
Since $\phi(I_{\Omega})=I_{\Omega}$ and $\phi(J_{\Omega})=J_{\Omega}$, the subspace $^{\Phi}\!\!{\mathcal{A}}=\{X\in{\mathcal{A}}\mid\forall_{\phi\in\Phi}\ \phi(X)=X\}$ is a coherent Jordan algebra. It is easy to verify that the matrices $\{\underline{\Phi(s)}\}_{s\in S}$ form the standard basis of $^{\Phi}\!\!{\mathcal{A}}$. ∎
It follows from above that given any $\Phi\leq\mathsf{JAut}({\mathfrak{S}})$, the pair $^{\Phi}{{\mathfrak{S}}}:=(\Omega,\{\Phi(s)\}_{s\in S})$ is also a Jordan configuration which we will refer to as an algebraic fusion of $\mathfrak{S}$}.
@@ 2.6. Constructing Jordan configurations from coherent ones
If ${\mathfrak{S}}=(\Omega,S)$ is a coherent configuration, then it is also a Jordan configuration. Recall that a permutation $\phi$ of $S$ is called an algebraic automorphism (anti-automorphism) of $\mathfrak{S}$} if its linear extension to $\mathcal{A}:=\langle\underline{s}\rangle_{s\in S}$ is an automorphism (anti-automorphism) of $(\mathcal{A},\cdot)$. The group $\mathsf{AAut}({\mathfrak{S}})$ of all algebraic automorphisms of $\mathfrak{S}$} is contained in $\mathsf{JAut}({\mathfrak{S}})$. Also, $\tau\in\mathsf{JAut}({\mathfrak{S}})$. Every anti-automorphism of $\mathfrak{S}$} is a composition of $\tau$ and some $\phi\in\mathsf{AAut}({\mathfrak{S}})$. All automorphisms and anti-automorphisms of $\mathfrak{S}$} form a subgroup of $\mathsf{JAut}({\mathfrak{S}})$ which will be denoted by $\mathsf{TAut}({\mathfrak{S}})$. If $\mathfrak{S}$} is commutative, then $\mathsf{AAut}({\mathfrak{S}})=\mathsf{TAut}({\mathfrak{S}})=\mathsf{JAut}({\mathfrak{S}})$. If however $\mathfrak{S}$} is not commutative, then $\tau\not\in\mathsf{AAut}({\mathfrak{S}})$ and $\mathsf{TAut}({\mathfrak{S}}):\mathsf{AAut}({\mathfrak{S}})]=2$. It is easy to check that $\mathsf{TAut}({\mathfrak{S}})=\mathsf{AAut}({\mathfrak{S}})\times\langle\tau\rangle$. In general, the inclusion $\mathsf{TAut}({\mathfrak{S}})\leq\mathsf{JAut}({\mathfrak{S}})$ may be proper. The smallest examples are $\mathrm{CC}(8,31)$ and $\mathrm{CC}(8,32)$ of order $8$. (Enumeration of coherent configurations here is taken from Ziv-Av’s catalog [20] of small coherent configurations.
If ${\mathfrak{S}}=(\Omega,S)$ is a thin association scheme, then $S$ is a regular subgroup of ${\sf Sym}(\Omega)$. In this case a permutation $\phi\in{\sf Sym}(S)$ belongs to $\mathsf{JAut}({\mathfrak{S}})$ if and only if it satisfies $\phi(st)\in\{\phi(s)\phi(t),\phi(t)\phi(s)\}$ for all $s,t\in S$. It was shown in [15] that any such map is either an automorphism or anti-automorphism111 The authors are grateful to R. Lipyanski who informed them of the paper [15].. Thus for thin association schemes one always has $\mathsf{TAut}({\mathfrak{S}})=\mathsf{JAut}({\mathfrak{S}})$.
We call a Jordan configuration $\mathfrak{S}=(\Omega,S)$ non-autonomous if it is an algebraic fusion of a coherent configuration $\mathfrak{C}=(\Omega,C)$ via a subgroup $\Phi\leq\mathsf{JAut}(\mathfrak{C})$. A Jordan configuration will be called autonomous otherwise, that is, if it cannot be obtained as an algebraic fusion of some coherent configuration sharing the same point set.
Note that in [14] Jordan schemes were divided into two types: proper and improper. An improper Jordan scheme is one that can be realized as a symmetrization of an association scheme. Otherwise it is proper. Under this definition, any nonsymmetric Jordan scheme is automatically proper. Regarding symmetric Jordan schemes, there are proper ones that are non-autonomous. We are able to show that there are no examples of proper non-autonomous Jordan schemes of order less than $16$. In fact, the smallest examples of such schemes known to us are of order $24$ as described below.
@@
Computer calculations show that each of $\text{AS}(24,526)$, $\text{AS}(24,597)$, $\text{AS}(24,669)$ has two algebraic fusions that are proper symmetric Jordan schemes. Hence, it follows that these schemes are non-autonomous. Enumeration of association schemes here is taken from Hanaki and Miyamoto’s catalog [7] of small association schemes. These are also accessible in GAP [6] as well as in the COCO2P package [12].
It appears that the proper Jordan schemes constructed in [14] are also autonomous, but we have yet to prove this. We have however verified this for the smallest example of order $15$ constructed there.
@@ 3. Thin Jordan schemes
It is well known that each basic graph of an association scheme $(\Omega,S)$ is regular. Therefore we always have that $|S|\leq|\Omega|$ where equality holds if and only if each basic relation $s\in S$ is a permutation on $\Omega$. Such a scheme is called thin and it is not difficult to check that an association scheme $(\Omega,S)$ is thin if and only if $S$ is a regular subgroup of ${\sf Sym}(\Omega)$. Since every group has a regular representation, one can consider groups as a particular case of association schemes.
Although symmetric Jordan schemes are always regular, this is no longer true of nonsymmetric ones as the following simple example illustrates.
@@
The full matrix algebra $M_{\Omega}({\mathbb{F}})$ is a discrete coherent algebra of rank $|\Omega|^{2}$. In the case of $\Omega=\{1,2\}$ the mapping $\phi:\begin{bmatrix}a&b\\ c&d\\ \end{bmatrix}\rightarrow\begin{bmatrix}d&b\\ c&a\\ \end{bmatrix}$ is an anti-automorphism. Therefore, the subspace ${\mathcal{A}}:={}^{\!\langle\phi\rangle\!}M_{\{1,2\}}({\mathbb{F}})=\left\{\begin{bmatrix}a&b\\ c&a\\ \end{bmatrix}\,\vline\,a,b,c\in{\mathbb{F}}\right\}$ is a coherent J-algebra which has order 2 and rank 3. The corresponding Jordan configuration $(\Omega,\{s_{0},s_{1},s_{2}\})$, where $s_{0}=I_{\Omega}$, $s_{1}=\{(1,2)\}$ and $s_{2}=\{(2,1)\}$, is homogeneous but not regular, and its rank-to-order ratio $|S|/|\Omega|$ is $3/2$.
The following assertion shows that this is the maximum value of a rank-to-order ratio for any nonregular Jordan scheme.
@@
Let $(\Omega,S)$ be a nonregular Jordan scheme. Then $|S|/|\Omega|\leq 3/2$ where equality holds if and only if all basic relations are thin.
@@
By [14, Proposition 2.2] there exists a partition $\Omega=\Omega_{1}\cup\Omega_{0}$ with $|\Omega_{1}|=|\Omega_{0}|$ such that every regular relation is contained in $\Omega_{0}^{2}\cup\Omega_{1}^{2}$ and each nonregular one is contained in either $\Omega_{0}\times\Omega_{1}$ or $\Omega_{1}\times\Omega_{0}$. Thus $S$ splits into a disjoint union $S=S_{r}\cup S_{0,1}\cup S_{1,0}$ where $S_{r}$ stands for all regular relations while $S_{i,1-i}:=\{s\in S\mid\subseteq\Omega_{i}\times\Omega_{1-i}\}$.
Define $n_{s}:=\frac{1}{2}p_{\{s,s^{t}\}}^{1_{\Omega}}$, for each $s\in S_{i,1-i}$. Then $|s|=n_{s}|\Omega_{i}|$, and it follows from $\sum_{s\in S_{i,1-i}}|s|=|\Omega_{i}|\cdot|\Omega_{1-i}|$ that $\sum_{s\in S_{i,1-i}}n_{s}=|\Omega|/2$. Since all summands in the latter sum are nonnegative integers, we conclude $|S_{i,1-i}|\leq|\Omega|/2$ for $i=0,1$.
@@
Arguing as before we conclude that $|S_{r}|\leq|\Omega|/2$. Combining together all of the above, we obtain
@@
Moreover, by the above arguments it follows that $|S|=\frac{3}{2}|\Omega|$ if and only if $n_{s}=1$ for all $s\in S$. ∎
Since the maximum rank-order ratio is attained by nonregular thin Jordan schemes, our next goal is to classify them. This will be proceeded in the next subsection.
3.1. Nonregular thin Jordan schemes
Let us begin by describing how every commutative homogeneous coherent algebra gives rise to a nonregular homogeneous coherent Jordan algebra.
@@ 3.1. Nonregular thin Jordan schemes
Let $\mathcal{C}\leq M_{n}(\mathbb{F})$ be a commutative homogeneous coherent algebra. We define $\mathcal{J}(\mathcal{C})\subseteq M_{2n}(\mathbb{F})$ as follows:
@@
Clearly, $\mathcal{J}(\mathcal{C})$ is a linear subspace of $M_{2n}(\mathbb{F})$. Since $\mathcal{C}$ is itself closed under Schur-Hadamard multiplication, so is $\mathcal{J}(\mathcal{C})$.
It remains to show closure of the Jordan product. To this end, let $A,B,C,D,E,F\in\mathcal{C}$. Then
@@
An important special case is when $\mathcal{C}$ is the adjacency algebra of a thin association scheme coming from an abelian group $G$. In this case, we shall write $\mathcal{J}(G)$ rather than the more accurate, albeit cumbersome, $\mathcal{J}(\mathcal{A}(G,G))$. Clearly, for every abelian group $G$ we have that $\mathcal{J}(G)$ is a nonregular thin homogeneous coherent Jordan algebra.
@@
Let $\mathfrak{S}=(\Omega,\mathcal{C})$ be a nonregular thin Jordan scheme. and let $\mathcal{A}$ be its adjacency algebra. Then there exists an abelian group $G$ and a permutation matrix $P$, such that $\mathcal{A}=P^{-1}\mathcal{J}(G)P$.
@@
Let $\mathfrak{J}=(\Omega,\mathcal{C})$ be a nonregular thin Jordan scheme and let $\mathcal{A}$ be its adjacency algebra. Let ${\mathcal{S}}=\{A_{i}\}_{i=1}^{r}$ be the standard basis of $\mathcal{A}$. By [14, Proposition 2.2], each element of $\mathcal{S}$} is of one of the following shapes:
@@
where $A,B,C,D$ are certain matrices of order $n$, and where $O$ denotes the zero matrix. Since $\mathfrak{S}$ is thin, each of $A,B,C,D$ is a permutation matrix.
Let us choose the following designations:
@@
Note that $\mathcal{K}=\mathcal{H}^{t}=\mathcal{H}^{-1}$ and $\sum\mathcal{G}_{1}=\sum\mathcal{G}_{2}=\sum\mathcal{H}=\sum\mathcal{K}=J$. Moreover, $|\mathcal{G}_{1}|=|\mathcal{G}_{2}|=|\mathcal{H}|=|\mathcal{K}|=n$.
We next prove
@@
Suppose $\left[\begin{smallmatrix}A&O\\[2.84526pt] O&B_{1}\end{smallmatrix}\right],\left[\begin{smallmatrix}A&O\\[2.84526pt] O&B_{2}\end{smallmatrix}\right]\in{\mathcal{S}}$ with $B_{1}\neq B_{2}$. Then $\left[\begin{smallmatrix}O&O\\[2.84526pt] O&B_{1}-B_{2}\end{smallmatrix}\right]\in\mathcal{A}$, whence also
$\begin{bmatrix}O&O\\[2.84526pt] O&B_{1}-B_{2}\end{bmatrix}\circ\begin{bmatrix}A&O\\[2.84526pt] O&B_{1}\end{bmatrix}=\begin{bmatrix}O&O\\[2.84526pt] O&B_{1}\end{bmatrix}\in{\mathcal{A}}.$
But since $\mathcal{A}$ is closed under taking powers, we thereby obtain $\left[\begin{smallmatrix}O&O\\[2.84526pt] O&I\end{smallmatrix}\right]\in\mathcal{A}$, a contradiction under the assumption that $\mathfrak{J}$ is a Jordan-scheme. It therefore follows that $B_{1}=B_{2}$.
This establishes a bijection between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ which we denote by $A\mapsto\varphi(A)$. Clearly, $\mathcal{G}_{1}$ is the standard basis of a thin homogeneous coherent $J$-algebra, and likewise for $\mathcal{G}_{2}$.
We next show that
@@
To see this, first observe that for all $C\in\mathcal{H}$, $D\in\mathcal{K}$ we have
@@
Since $|C\mathcal{K}|=|\mathcal{K}|=|\mathcal{G}_{1}|$, there exists $D\in\mathcal{K}$ for which $A=CD$. By (1), $\varphi(A)=DC$ holds for this $D$.
Next we establish
@@
To this end, let $A\in\mathcal{G}_{1}$. By (2), there exists $D\in\mathcal{K}$ such that $A=CD$ and $\varphi(A)=DC$. From this it follows that $D=C^{-1}A$, hence $\varphi(A)=C^{-1}AC$.
Let $P$ be a permutation matrix and let $X\in{\mathcal{A}}$ be arbitrary. Then conjugation $X\mapsto PXP^{-1}$ preserves each of the operations $\cdot$, $\tau$, $\star$, $\circ$ on $\mathcal{A}$}. In other words, the resulting algebra $\mathcal{A}^{\prime}=P\mathcal{A}P^{-1}$ is a homogeneous coherent $J$-algebra and their corresponding Jordan schemes are combinatorially isomorphic.